# Bandit Sequential Posted Pricing via Half-Concavity

### v1

@@ 4 $n$ Buyers with General Distributions
For BSPP problem with $n$ buyers, there exists an algorithm with $O(n^{5/3}T^{2/3}\log T)$ regret.
4.1 Proof Overview
Comparing to the regular buyers setting, one extra challenge for general buyers setting is that the revenue function in this case is no longer single-peakedness, leading to the failure of the idea of confidence intervals. Therefore, the first step of our algorithm is to discretize the value space: it can be shown that rounding buyers’ valuations to multiples of $\epsilon$ only brings an extra $\epsilon$ loss. The idea of discretization was first used in [KL03] for single buyer setting. We will show that the same idea also works for sequential buyers.
After discretizing the value space, it suffices to solve BSPP for general distributions with discrete support. Specifically, we have the following theorem:
Theorem 4.2.
The main structure of the algorithm given by 4.2 is almost identical to the algorithm for regular buyers - we design a subroutine that shrinks the possible candidates of $p^{*}_{i}$ while keeping a low regret inside the new set of candidates. Calling this subroutine $O(\log T)$ times with a halving error parameter gives the desired algorithm. Comparing to the regular distributions setting, the idea of confidence intervals does not work for discrete distributions, because the candidates may be discontinuous. To fix this issue, we use $n$ price sets to maintain possible candidate prices.
@@ 4.2 BSPP is Discretizable
Proof of 4.3.
Let $p^{*}_{1},\cdots,p^{*}_{n}$ be the optimal prices for $R(p_{1},\cdots,p_{n})$, and let $p^{\prime}_{i}=\lfloor\frac{p^{*}_{i}}{\epsilon}\rfloor\cdot\epsilon$. To prove 4.3, we only need to show
@@
With the help of 4.3, a corollary of 4.2 directly leads to 4.1:
@@
We take $k=n^{-5/3}T^{1/3}$ and uniformly discretize $[0,1]$ into $k$ prices. 4.3 suggests that this leads to an extra $\frac{1}{k}=n^{5/3}T^{-1/3}$ error, so the total regret from the discretization is $T\cdot\frac{1}{k}=O(n^{5/3}T^{2/3})$. On the other hand, the total regret from the algorithm given by 4.2 is $O(n^{2.5}\sqrt{kT}\log T+n^{5}k\log^{2}T)=O(n^{5/3}T^{2/3}\log T+n^{10/3}T^{1/3}\log^{2}T)$. In this paper, we further assume that $n^{5/3}\log T=o(T^{1/3})$, i.e., $n^{5}\log^{3}T=o(T)$, otherwise an $\widetilde{O}(\mathsf{poly}(n))$ algorithm is trivial. Then, the total regret is $O(n^{5/3}T^{2/3}\log T)$. ∎
4.3 Algorithm for Discrete Buyers: Proof of 4.2
In this section, we finish our proof of 4.1 by presenting the algorithm given by 4.2.
@@ 4.3.1 Notation
We use the same set of the notations as the regular distributions algorithm. Specifically, we will use the following notations:
@@
  • $\mathsf{CP}_{i}$: The candidate prices set of buyer $i$. Initially $\mathsf{CP}_{i}$ is set to be $V$, and it will be gradually shrinked during the algorithm.
@@
  • $R_{i}(p)$: We define $R_{i}(p)~{}:=~{}R\big{(}r_{1},\cdot,r_{i-1},p_{i}=p,\hat{p}_{i+1},\cdots,\hat{p}_{n}\big{)}.$ It should be noted that $R_{i}(p)$ can be also written as
    $R_{i}(p)~{}=~{}\mathsf{Pre}(S_{i})+P_{i}\big{(}p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p)\big{)}.$
@@ 4.3.2 High-Level Structure: Shrinking Candidate Prices Sets
We defer the detailed algorithm and proofs for 4.4 to Section 4.3.3 and first give the proof for 4.2. The main idea is to use 4.4 for $O(\log T)$ times and gradually halve the one-round error when playing with prices inside the candidate sets.
@@ Acknowledgements
We are thankful to Thomas Kesselheim as the project was initiated in discussions with him. We also thank the anonymous reviewers who helped improve the presentation of the paper.
Appendix A Missing Proofs from Section 2
A.1 An Alternate Proof of Half-Concavity via Concavity in Quantile Space
Proof of the Half-Concavity in 2.4.
We prove by contradiction. Assume there exist $0\leq a<b<c\leq p^{*}$ satisfying $c=b+t(b-a)$ and $R(c)>R(b)+t\cdot\big{(}R(b)-R(a)\big{)}$.
$R(c)$ $~{}>~{}R(b)+t\cdot(R(b)-R(a))$
$~{}=~{}b\cdot q(b)+t\cdot\big{(}b\cdot q(b)-a\cdot q(a)\big{)}$
$~{}=~{}\big{(}(t+1)b-ta\big{)}\cdot\big{(}(t+1)q(b)-tq(a)\big{)}$
$~{}=~{}c\cdot\overline{q}~{}\geq~{}c\cdot q(c)~{}=~{}R(c).\qed$
A.2 Proof of 2.9
See 2.9
We first give the following Algorithm 9, and then show Algorithm 9 is the desired algorithm for 2.9.
1 Let $\delta=\frac{\epsilon}{100}$ be the scaled error parameter.
// First binary search to determine $\ell^{\prime}$
3 Let $\ell_{b1}\leftarrow\ell,r_{b1}\leftarrow\hat{p}$.
4 while $r_{b1}-\ell_{b1}\geq{T^{-100}}$ do
5       Let $m=\frac{\ell_{b1}+r_{b1}}{2}$.
8
9Let $\ell^{\prime}\leftarrow\ell_{b1}$
// Second binary search to determine $r^{\prime}$
11 if $\hat{R}(r-{T^{-100}})\geq\hat{R}(\hat{p})-2\delta$ then
       // Special Case: original upper bound is good
12       Let $r^{\prime}\leftarrow r$
13else
14       Let $\ell_{b2}\leftarrow\hat{p},r_{b2}\leftarrow r-{T^{-100}}$.
15       while $r_{b2}-\ell_{b2}\geq{T^{-100}}$ do
16             Let $m=\frac{\ell_{b2}+r_{b2}}{2}$.
19
20      Let $r^{\prime}\leftarrow r_{b2}$.
Output: $[\ell^{\prime},r^{\prime}]$.
ALGORITHM 9 Getting $[\ell^{\prime},r^{\prime}]$ via Binary Search
Proof of 2.9.
Claim A.1.
Proof.
We will give the remaining parts of the proof assuming that A.1 holds.
Now we show $p^{*}\in[\ell^{\prime},r^{\prime}]$. We first prove $p^{*}\geq\ell^{\prime}$. In Algorithm 9, the only way we update $\ell^{\prime}$ is that we observe $\hat{R}(m)<\hat{R}(\hat{p})-2\delta$ in the first binary search, and $\ell^{\prime}$ is updated to at least $m$. This is feasible because A.1 gives $R(m)<R(\hat{p})\leq R(p^{*})$, and the single-peakedness of $R(p)$ guarantees $m\leq p^{*}$. The proof of $p^{*}\leq r^{\prime}$ is symmetric.
We first prove that $R(\hat{p})-R(\ell^{\prime})\leq 5\delta$. The binary search subroutine guarantees that $\hat{R}(r_{b1})\geq\hat{R}(\hat{p})-2\delta$ after the binary search loop is finished. Then, A.1 ensures that $R(\hat{p})-R(r_{b1})\leq 4\delta$. Since $\ell^{\prime}$ is set to be $\ell_{b1}>r_{b1}-{T^{-100}}$, the 1-Lipschitzness of $R(p)$ in $[\ell_{b1},\hat{p}]$ guarantees that
$R(r^{\prime}-{T^{-100}})~{}\geq~{}R(\ell_{b2})~{}\geq~{}R(\hat{p})-4\delta.$
Otherwise, the 1-Lipschitzness gives
Appendix B Missing Proof from Section 3
B.1 Proof of Partial Half-Concavity of $R_{i}(p)$
See 3.6
Proof.
[…]
This means $R_{i}(\mathsf{Rev}(S_{i+1}))$ is no worse than any price in $[0,\mathsf{Rev}(S_{i+1})]$. Therefore, there must be $\arg\max_{p}R_{i}(p)\geq\mathsf{Rev}(S_{i+1})$.
$\frac{R(c)}{P_{i}}~{}>~{}$ $\frac{R(b)}{P_{i}}+\frac{t\cdot(R(b)-R(a))}{P_{i}}$
[…]
$-t(t+1)(b-a)(q(a)-q(b))$
[…]
B.2 Proof of 3.9
See 3.9
Proof.
where the last inequality holds when $C$ is a constant greater than $60$.
[…]
where the last inequality holds when $C$ is a constant greater than $16$. ∎
B.3 Proof of 3.12: Generalized Algorithm for Half $\lambda$-Concave Functions
In this section, we give the generalized algorithm for half $\lambda$-concave functions. The definition of half $\lambda$-concave functions and the main lemma for the algorithm are stated as follows:
See 3.10
See 3.12
Algorithm 10 is the desired algorithm for 3.12. The algorithm is almost the same as Algorithm 2. The only difference is that extra care is needed for the extra parameter $\lambda$.
1 Fix $p_{j}=r^{s}_{j}$ for $j<i$ and $p_{j}=\hat{p}_{j}$ for $j>i$.
3 while $r_{j}-\ell_{j}>\delta$ do
7       if $\hat{R}_{i}(a_{j})<\hat{R}_{i}(b_{j})-2\delta$ then
8            Let $\ell_{j+1}\leftarrow a_{j},r_{j+1}\leftarrow r_{j}$
9      else
10            Let $\ell_{j+1}\leftarrow\ell_{j},r_{j+1}\leftarrow b_{j}$
11      $j\leftarrow j+1$
Output: $\hat{p}_{i}$.
Proof of 3.12.
Claim B.1.
Proof.
Now, we prove $R_{i}(p^{*}_{\ell,r})-R_{i}(\hat{p}_{i})\leq 6\lambda\delta$ assuming B.1 holds. We demonstrate the proof by discussing the following two cases. The first case is that there exists $j\in[k]$, such that $p^{*}_{\ell,r}$ falls in $[\ell_{j},r_{j}]$ but not in $[\ell_{j+1},r_{j+1}]$. In this case, there must be $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ and $[b_{j},r_{j}]$ is dropped. If not, we have $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ and $[\ell_{j},a_{j}]$ is dropped. However, the inequality $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ together with B.1 gives $R(a_{j})<R(b_{j})$. Then, the single-peakedness of $R_{i}(p)$ gives $p^{*}_{\ell,r}\geq a_{j}$, which is in contrast to the assumption that $p^{*}_{\ell,r}\notin[\ell_{j+1},r_{j+1}]$. Therefore, we have $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$, and the assumption $p^{*}_{\ell,r}\notin[\ell_{j+1},r_{j+1}]$ implies $p^{*}_{\ell,r}\in[b_{j},r_{j}]$. Then, the half $\lambda$-concavity of $R_{i}(p)$ in $[a_{j},p^{*}_{\ell,r}]$ guarantees
B.4 Proof of 3.14
See 3.14
We first give the following Algorithm 11, and then show Algorithm 11 is the desired algorithm for 3.14.
2 Let $\ell_{b1}\leftarrow\ell_{i},r_{b1}\leftarrow\hat{p}_{i}$.
3 while $r_{b1}-\ell_{b1}\geq{T^{-100}}$ do
4       Let $m=\frac{\ell_{b1}+r_{b1}}{2}$.
7
  // Second binary search to determine $r^{\prime}_{i}$
11 if $\hat{R}_{i}(r-{T^{-100}})\geq\hat{R}_{i}(\hat{p}_{i})-2\delta$ then
       // Special Case: original upper bound is good
12       Let $r^{\prime}_{i}\leftarrow r_{i}$
13else
14       Let $\ell_{b2}\leftarrow\hat{p}_{i},r_{b2}\leftarrow r_{i}-{T^{-100}}$.
15       while $r_{b2}-\ell_{b2}\geq{T^{-100}}$ do
16             Let $m=\frac{\ell_{b2}+r_{b2}}{2}$.
19
20      Let $r^{\prime}_{i}\leftarrow r_{b2}$.
Output: $[\ell^{\prime}_{i},r^{\prime}_{i}]$.
ALGORITHM 11 Getting $[\ell^{\prime}_{i},r^{\prime}_{i}]$ via Binary Search
Proof of 2.9.
We show Algorithm 11 satisfies the statements in 3.14. Our proof starts from verifying Algorithm 11 runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds with $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$. For the number of tested rounds, Algorithm 11 runs two standard binary searches. In one binary search, the subroutine tests one price with $C\cdot\frac{\log T}{\delta^{2}}$ rounds, and the binary search stops when the length of the interval is less than ${T^{-100}}$. Therefore, The total number of rounds is $C\cdot\frac{\log T}{\delta^{2}}\cdot O(\log T^{100})=O(\frac{\log^{2}T}{\epsilon^{2}})$. For the range constraint of all tested prices, Algorithm 11 directly guarantees that every tested price falls in $[\ell_{i},r_{i}-{T^{-100}}]$.
Claim B.2.
Proof.
Now we show $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$. The key idea is to use3.13. For clarity we first restate 3.13: See 3.13
Now we prove $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$. We first show that $p^{*}_{i}\geq\ell_{b1}$. Here, $\ell_{b1}$ refers to the value after the while loop is finished. $p^{*}_{i}\geq\ell_{b1}$ holds because $\ell_{b1}$ is updated to be at least $m$ when we observe $\hat{R}_{i}(m)<\hat{R}_{i}(\hat{p}_{i})-2\delta-5(n-i)\delta$. Combining this with B.2 implies $R_{i}(m)<R_{i}(\hat{p}_{i})-5(n-i)\delta$, i.e., $m\notin I^{\prime}$. Since $R_{i}(p)$ is single-peaked, it’s sufficient to update $\ell_{b1}$ to at least $m$ and $I^{\prime}\subseteq[\ell_{b1},1]$ holds. Symmetrically, a similar argument gives $p^{*}_{i}\leq r_{b2}=r^{\prime}_{i}$.
It remains to show $p^{*}_{i}\geq\ell^{\prime}_{i}$. $\ell^{\prime}_{i}$ and $\ell_{b1}$ are different when we have $\hat{R}(\ell_{b1})<\hat{R}(r_{b1})-3\delta$. Combining this inequality with B.2 gives $R(r_{b1})-R(\ell_{b1})>\delta>r_{b1}-\ell_{b1}$. This is in contrast to the constrained Lipschitzness from 3.6. Then, there must be $r_{b1}<\mathsf{Rev}(S_{i+1})<\mathsf{Opt}(S_{i+1})\leq p^{*}_{i}$, where the last inequality is from 3.5. Therefore, it’s feasible to update $\ell^{\prime}_{i}$ to $r_{b1}$ in this case.
We first prove that $R_{i}(\hat{p}_{i})-R(\ell^{\prime}_{i})\leq 5(n-i)\delta+11\delta$. The binary search subroutine guarantees that $\hat{R}_{i}(r_{b1})\geq\hat{R}_{i}(\hat{p}_{i})-2\delta-5(n-i)\delta$ after the binary search loop is finished. Then, B.2 ensures that $R_{i}(\hat{p}_{i})-R_{i}(r_{b1})\leq 5(n-i)\delta+4\delta$. Then, if $\hat{R}_{i}(\ell_{b1})<\hat{R}_{i}(r_{b1})-3\delta$, we set $\ell^{\prime}_{i}$ to be $r_{b1}$. Otherwise, we have $\hat{R}_{i}(\ell_{b1})\geq\hat{R}_{i}(r_{b1})-3\delta$ and $\ell^{\prime}_{i}$ is set to be $\ell_{b1}$. Then,
[…]
Otherwise, the constrained Lipschitzness of $R_{i}(p)$ gives

### v2

@@ 4 $n$ Buyers with General Distributions
For BSPP problem with $n$ buyers, there is an algorithm with $O(n^{5/3}T^{2/3}\log T)$ regret.
Algorithm Overview.
Comparing to the regular buyers setting, one extra challenge for general buyers setting is that the revenue function in this case is no longer single-peaked, leading to the failure of the idea of confidence intervals. Therefore, the first step of our algorithm is to discretize the value space: it can be shown that rounding buyers’ valuations to multiples of $\epsilon$ only brings an extra $\epsilon$ loss. The idea of discretization was first used in [KL03] for single buyer setting. We will show that the same idea also works for sequential buyers. In this section, we take the discretization parameter $\epsilon=n^{5/3}T^{-1/3}$, and the accumulated regret from the discretization step would be $T\cdot O(\epsilon)=O(n^{5/3}T^{2/3})$.
After discretizing the value space, it suffices to solve BSPP for general distributions with discrete support. The main idea of our algorithm is almost identical to the algorithm for regular buyers - we design a subroutine that shrinks the possible candidates of $p^{*}_{i}$ while keeping a low regret inside the new set of candidates. Calling this subroutine $O(\log T)$ times with a halving error parameter gives the desired algorithm. Comparing to the regular distributions setting, the idea of confidence intervals does not work for discrete distributions, because the candidates may be discontinuous. To fix this issue, we use $n$ price sets to maintain possible candidate prices.
@@ 4.1 BSPP is Discretizable
Proof of 4.2.
Let $p^{*}_{1},\cdots,p^{*}_{n}$ be the optimal prices for $R(p_{1},\cdots,p_{n})$, and let $p^{\prime}_{i}=\lfloor\frac{p^{*}_{i}}{\epsilon}\rfloor\cdot\epsilon$. To prove 4.2, we only need to show
@@
4.2 Algorithm for Discrete Buyers
Now we give our algorithm for buyers with discrete support. Our main goal is to prove the following 4.3:
Theorem 4.3.
With the help of 4.3, combining 4.2 with 4.3 proves the main4.1. Therefore, proving 4.3 is sufficient to give the desired $T^{2/3}$ regret bound for general buyers.
@@ 4.2 Algorithm for Discrete Buyers
We take $k=n^{-5/3}T^{1/3}$ and uniformly discretize $[0,1]$ into $k$ prices. 4.2 suggests that this leads to an extra $\frac{1}{k}=n^{5/3}T^{-1/3}$ error, so the total regret from the discretization is $T\cdot\frac{1}{k}=O(n^{5/3}T^{2/3})$. On the other hand, the total regret from the algorithm given by 4.3 is $O(n^{2.5}\sqrt{kT}\log T+n^{5}k\log^{2}T)=O(n^{5/3}T^{2/3}\log T+n^{10/3}T^{1/3}\log^{2}T)$. In this paper, we further assume that $n^{5/3}\log T=o(T^{1/3})$, i.e., $n^{5}\log^{3}T=o(T)$, otherwise an $\widetilde{O}(\mathsf{poly}(n))$ algorithm is trivial. Then, the total regret is $O(n^{5/3}T^{2/3}\log T)$. ∎
@@ 4.2.1 Notation
Our proof of 4.3 starts from introducing the notations. We use the same set of the notations as the regular distributions algorithm, which are restated for convenience:
@@
  • $R_{i}(p)$: We define $R_{i}(p)~{}:=~{}R\big{(}r_{1},\cdot,r_{i-1},p_{i}=p,\hat{p}_{i+1},\cdots,\hat{p}_{n}\big{)}.$ It should be noted that $R_{i}(p)$ can be also written as $R_{i}(p)~{}=~{}\mathsf{Pre}(S_{i})+P_{i}\big{(}p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p)\big{)}.$
@@
We also introduce the following notations for the discretized setting:
  • $\mathsf{CP}_{i}$: The candidate prices set of buyer $i$. Initially $\mathsf{CP}_{i}$ is set to be $V$.
@@ 4.2.2 High-Level Structure: Shrinking Candidate Prices Sets
We defer the detailed algorithm and proofs for 4.4 to Section 4.2.3 and first give the proof for 4.3. The main idea is to use 4.4 for $O(\log T)$ times and gradually halve the one-round error when playing with prices inside the candidate sets.
@@ Acknowledgements
We are thankful to Thomas Kesselheim as the project was initiated in discussions with him. We also thank the anonymous reviewers who helped greatly improve the presentation of the paper.